OpenStudy (zenmo):

Implicit differentiation part 2

OpenStudy (zenmo):

\[9x^2+y^2=9\]

OpenStudy (zenmo):

What I have so far: \[18x+2yy'=0\] \[2yy'=-18x\] \[y'=\frac{ -9x }{ y }\] Using quotient rule: \[y'' = - 9 \times \frac{ y \times 1 - x \times y' }{ y^2 }=-9\times \frac{ y-x(\frac{ -9x }{ y }) }{ y^2 }\] \[= - 9 \times \frac{ y^2+9x^2 }{ x^3 }\]

OpenStudy (zenmo):

Now, I just need to convert that into an answer.

zepdrix (zepdrix):

Oh second derivative? Interestinggg :)

OpenStudy (zenmo):

yep

zepdrix (zepdrix):

Looks good up to this point,\[y'' = - 9 \times \frac{ y \times 1 - x \times y' }{ y^2 }=-9\times \frac{ y-x(\frac{ -9x }{ y }) }{ y^2 }\]I'm trying to figure out what happened on the next line.. hmm

zepdrix (zepdrix):

Oh the denominator is y^3, small typo, ok great

OpenStudy (zenmo):

opps, yea it is y^3

OpenStudy (zenmo):

Do you know how to simplify that further?

OpenStudy (zenmo):

the book does it as: \[-9 \times \frac{ 9 }{ y^3 }\]

zepdrix (zepdrix):

Ooo you see it, don't you? hehe

zepdrix (zepdrix):

\[\large\rm y''=-9\frac{\color{orangered}{y^2+9x^2}}{y^3}\]And our original expression is given to be\[\large\rm \color{orangered}{9x^2+y^2=9}\]

zepdrix (zepdrix):

You didn't reduce the 9 correctly, don't worry about needing to do that anyway.

OpenStudy (zenmo):

oh i see it now

zepdrix (zepdrix):

You're just replacing the numerator by 9, because that's what it is equivalent to.

OpenStudy (zenmo):

\[y^2+9x^2=9\]

zepdrix (zepdrix):

Just a not on the reduction though,\[\large\rm 9x^2+y^2=9\qquad\to\qquad x^2+\frac19y^2=1\]It doesn't quite work out the way you would like.

zepdrix (zepdrix):

a note*

OpenStudy (zenmo):

Okay, I understand it fully now

OpenStudy (zenmo):

I figured I had to simplify that part, but didn't see that the last piece was the original equation itself y^2+9x^2 = 9

zepdrix (zepdrix):

yay good job

OpenStudy (zenmo):

thanks :)